A container has 2.4 liters of water at 20^C.    What is the amount of heat required to boil the water?  Write final answer to the nearest whole number and indicate the unit in kJ. 


1
Expert's answer
2021-10-24T01:28:27-0400

2.4l= 2400ml

density= 1ml/g

Mass= 1× 2400= 2400g= 2.4kg

Water boils at 0°C

Boiling point = 100°C

Initial = 20°C

Specific heat capacity of water = 4200j/kg/°C

= 4200×(100-20)×2.4

= 806400j

= 806.400kJ